1. Weston Corporation manufactures a product that is available in both a deluxe and a regular model. The company has made the regular model for years; the deluxe model was introduced several years ago to tap a new segment of the market. Since introduction of the deluxe model, the company’s profits have steadily declined. Sales of the deluxe model have been increasing rapidly. Overhead is applied to products on the basis of direct labor-hours. At the beginning of the current year, management estimated that $3,080,000 in overhead costs would be incurred and the company would produce and sell 10,000 units of the deluxe model and 50,000 units of the regular model. The deluxe model requires 2.0 hours of direct labor time per unit, and the regular model requires 1.0 hours. Materials and labor costs per unit are given below: Deluxe Regular Direct materials cost per unit $50.00 $30.00 Direct labor cost per unit $30.00 $15.00


1
Expert's answer
2021-09-24T06:16:44-0400

Total Budgeted Labor hours = 8,500*2 + 42,500*1.5 = 80,750 hours 1-a Predetermined rate = Budgeted Overheads/Budgeted Hours = 3,230,000/80,750